OpenStudy (anonymous):

How do you clear the equations of fractions?? E.G. Solve the Equation: x/2=10-x/3 PLEASE HELP!!!!!!! & Explain!

OpenStudy (anonymous):

6.(x/2)=6.(10-x/3)

OpenStudy (anonymous):

3.x=60-2.x

OpenStudy (anonymous):

6(x/2) =6(10-x/3) 3x = 60-2x 3x+2x=60 5x=60 x=60/5 x=12 :)

OpenStudy (anonymous):

5.x=60

OpenStudy (anonymous):

x=12

OpenStudy (anonymous):

Why is everyone multiplying by 6?

OpenStudy (anonymous):

to eliminate 3 and 2

OpenStudy (anonymous):

then why not use 3 itself and 2 itself? why 6?

OpenStudy (anonymous):

six is divisible for both

OpenStudy (anonymous):

o! so least common denominator(:

OpenStudy (anonymous):

thats it!

OpenStudy (anonymous):

yes lowest common multiple of 3 and 2 :)

OpenStudy (anonymous):

wait but in the middle of the problem, you guys moved 3x=60-2x to 3x+2x=60, why??

OpenStudy (anonymous):

you can just solve it like a linear equation and add 2x on both sides getting you 5x=60 which equals 12

OpenStudy (anonymous):

3x+2x = 60-2x+2x Eliminate the 2

OpenStudy (anonymous):

3.x=60-2.x->3.X+2.X=60-2.X+2X

OpenStudy (anonymous):

oh! alright, I see. Thanks.!(: